OpenStudy (anonymous):

solve for x 2/5x + 2 = 10

OpenStudy (anonymous):

Is this \[\frac{ x^2 }{ 5x+2 }=10\]

OpenStudy (anonymous):

no its 2 over 5 x +2

OpenStudy (anonymous):

\[\frac{ 2 }{ 5x +2 } = 10\]\[2 = 10 ( 5x + 2)\] \[2 = 50x + 20\]\[x = \frac{ -18 }{ 50 }\]\[x = \frac{ -9 }{ 25 }\]